OpenStudy (anonymous):

A.Find the distance between (-3,-5) and (9,0), the midpoint of the segment with these endpoints, then find the slope of the line containing them. B. Find the distance between(-2,-1) and (4,2), the midpoint of the segment with these endpoints, then find the slope of the line containing them.

OpenStudy (anonymous):

@Hero

OpenStudy (astrophysics):

Basically a bunch of formulas

OpenStudy (astrophysics):

\[\huge d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] distance formula \[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] slope formula \[\huge \left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]\] midpoint haha

OpenStudy (anonymous):

gotcha, all i need to do is plug it in. thanks!

OpenStudy (astrophysics):

Yup plug and chug!